# A new theory of continuum defects in terms of the general teleparallel geometry and exterior algebra

### v1

@@ Abstract
We discuss the geometric formulation of continuum defects consisting of dislocations and disclinations. After reviewing the metric affine geometry and the present geometric formulation of dislocation and disclination written in terms of torsion and full curvature (together with vanishing non-metricity), we give a new theory of them in a novel way in terms of torsion and non-metricity (together with vanishing full curvature), the so-called general teleparallel geometry. All calculations are performed by using the exterior algebra. We obtain continuity equations explicitly for dislocation density and disclination density.
@@ 1 Introduction
It is known that the role of geometry in physics is very valuable. The most familiar example of it is general relativity, Einstein’s theory of gravity, in which the gravitational field is associated with the Riemann curvature tensor. With similar motivations, geometric quantities have been used to express the theory of defects in materials. In studies conducted for this purpose so far, dislocation type defects have been paired with torsion and disclination type defects with full curvature [1],[2],[3],[4],[5],[6],[7],[8]. We review main results of those works in Section 3. There are works discussing the point defects apart from these two line defects as well [9],[10].
@@
However, if torsion is paired with dislocation and full curvature with disclination as in the geometric picture we summarized above, there will be no shift in dislocation when disclination occurs. On the contrary, the presence of dislocation causes a shift in disclination. This contradiction motivated us to do this work.
On the other hand, we know that a metric affine geometry is defined by three quantity; manifold, metric, affine connection. Furthermore, the affine connection 1-form, $\omega_{ab}$, contains a Riemannian component, the so-called Levi-Civita connection 1-form, $\widetilde{\omega}_{ab}$, computed from metric and a non-Riemannian component which we call as defect tensor-valued 1-form, $\mathrm{L}_{ab}$, consisting of tensors of torsion and non-metricity. Although in the literature of geometry and gravity it is usually called as distortion 1-form, we prefer to name it as defect 1-form appropriate to the topic. We summarize the basics of metric affine geometry in the forthcoming section. More specifically, while the symmetric part of defect 1-form consist of only non-metricity 1-form, its anti-symmetric part contains torsion and certain components of non-metricity. Thus, when the non-metricity vanishes, the defect 1-form turns out to be anti-symmetric and is given only by torsion. However, when non-metricity does not vanish, along with that the symmetric part of the defect 1-form appears, its anti-symmetric part is shifted by certain components of non-metricity. One can consult Ref.[18] for further reading. In conclusion, the resemblances between the statements on dislocation (disclination) and the anti-symmetric part (symmetric part) of defect 1-form led us to investigate the reformulation of the defect theory of a continuum in terms of Cartan-Weyl (general teleparallel) geometry. Besides all these, on the geometry side, the general teleparallel geometry has an advantage over the Riemann-Cartan geometry. While the latter does not have a metric formulation, the former does. In fact, all three teleparallel geometries have metric formulations [18]. As far as we know, our general teleparallel geometric formulation of continuum defects is the first attempt in the literature. Free from special discussion here, although plenty of papers has been published, no fundamental theory of defects is yet available. We develop our formulation for any non-Cartesian coordinates covariantly in the language of exterior algebra.
@@ 2 Metric affine geometry
The triple $\{M,g,\nabla\}$ defines a metric affine geometry where $M$ is three-dimensional orientable and differentiable manifold, $g$ is non-degenerate symmetric metric and $\nabla$ is full (or affine) connection or covariant derivative [19], [20]. We denote the anholonomic metric-orthonormal coframe by $e^{a}$, then write the metric as $g=g_{ab}e^{a}\otimes e^{b}$ where $g_{ab}:=\delta_{ab}$ is the Euclid metric with $\delta_{ab}=\text{diag}(1,1,1)$. For the holonomic coordinate coframe and the anholonomic mixed coframe we direct the reader to our paper [18]. The full connection is determined by the full connection 1-form $\omega^{a}{}_{b}$ via the definition $\nabla e^{a}:=-\omega^{a}{}_{b}\wedge e^{b}$ or $\nabla X_{a}:=X_{b}\,\omega^{b}{}_{a}$. An exterior product symbol $\wedge$ has not been inserted between $\omega^{b}{}_{a}$ and $X_{b}$, because the latter is not a differential form in the language of exterior algebra. It is supposed that there is a tensor product between them. Besides, $e^{a}$ is called the metric orthonormal (or shortly orthonormal) 1-form and the Cartan structure equations are given by tensor-valued non-metricity 1-form [^1] Here $\omega_{(ab)}=Q_{ab}$ is the reason why we adopt the factor $-\frac{1}{2}$ in the definition of non-metricity, tensor-valued torsion 2-form and tensor-valued full (or non-Riemannian) curvature 2-form, respectively,
$Q_{ab}$ $:=-\frac{1}{2}D\delta_{ab}=\omega_{(ab)},$ (1a)
@@
where $d$ is the exterior derivative and $D$ is the covariant exterior derivative. We utilize the usual notation: ()($\bullet$)( ∙ ) and []delimited-[][$\bullet$][ ∙ ] means to calculate the symmetric and anti-symmetric parts with respect to the indices enclosed, i.e., $\omega_{(ab)}=\frac{1}{2}(\omega_{ab}+\omega_{ba})$ and $\omega_{[ab]}=\frac{1}{2}(\omega_{ab}-\omega_{ba})$. They satisfy the Bianchi identities
@@
$D\mathfrak{T}^{a_{1}a_{2}\cdots a_{p}}_{\;\;\;\;b_{1}b_{2}\cdots b_{q}}=d\mathfrak{T}^{a_{1}a_{2}\cdots a_{p}}_{\;\;\;\;b_{1}b_{2}\cdots b_{q}}+\omega^{a_{1}}{}_{c}\wedge\mathfrak{T}^{ca_{2}\cdots a_{p}}_{\;\;\;\;b_{1}b_{2}\cdots b_{q}}+\cdots+\omega^{a_{p}}{}_{c}\wedge\mathfrak{T}^{a_{1}a_{2}\cdots c}_{\;\;\;\;b_{1}b_{2}\cdots b_{q}}$
$-\omega^{c}{}_{b_{1}}\wedge\mathfrak{T}^{a_{1}a_{2}\cdots a_{p}}_{\;\;\;\;cb_{2}\cdots b_{q}}-\cdots-\omega^{c}{}_{b_{q}}\wedge\mathfrak{T}^{a_{1}a_{2}\cdots a_{p}}_{\;\;\;\;b_{1}b_{2}\cdots c}.$ (3)
@@
Thus, we deduce that $Q_{abc}=Q_{(ab)c}$ is symmetric at the first two indices, $T^{a}{}_{bc}=T^{a}{}_{[bc]}$ is anti-symmetric at the last two indices and $R_{abcd}=R_{ab[cd]}$ is anti-symmetric at the last two indices and asymmetric $R_{abcd}=R_{[ab]cd}+R_{(ab)cd}$ at the first two indices by definition. Furthermore, it is worthy to remember that the symmetric part of the first two indices of $R_{abcd}$ depends on existence of the non-metricity and will allow us to define metrical disclination density tensor in the next section. Furthermore, we would like to remind explicitly that $d\delta_{ab}=0$, $d\delta^{ab}=0$ and $d\delta^{a}_{b}=0$, but $D\delta_{ab}=-2Q_{ab}$, $D\delta^{ab}=+2Q^{ab}$ and $D\delta^{a}_{b}=0$. Accordingly, one must pay special attention as raising or lowering an index in front of $D$ operation in non-metric geometries.
@@ 2.1 Decomposition of full connection
Here $\iota_{a}\equiv\iota_{X_{a}}$ denotes the interior product with respect to the orthonormal base vector $X_{a}$ such that $\iota_{b}e^{b}=\delta^{a}_{b}$. In the literature it is common to define the tensor-valued contortion 1-form, $K_{ab}=-K_{ba}$, in terms of torsion 2-form
@@ 2.2 Some algebraic identities
where * denotes the Hodge duality map, $Q=\delta_{ab}Q^{ab}$ is the non-metricity trace (or Weyl) 1-form, $\Lambda$ is any $p$-form, $\Gamma$ is any $q$-form and $pq$ is the ordinary multiplication of numbers of $p$ and $q$. The operation of Hodge map may be performed via the totally anti-symmetric epsilon tensor by relations
@@
2.3 Orthogonal decomposition of torsion and non-metricity
For the irreducible decomposition of torsion 2-form invariant under the orthogonal group $SO(3)$ we firstly define two suitable quantities; one-form $T$ and three-form $\sigma$
$T=\iota_{a}T^{a},\qquad\sigma=e_{a}\wedge T^{a}.$ (17)
Acoordingly, torsion can be split into three parts
$\underbrace{T^{a}}_{9}=\underbrace{\overset{(1)}{T}{}^{a}}_{5}+\underbrace{\overset{(2)}{T}{}^{a}}_{3}+\underbrace{\overset{(3)}{T}{}^{a}}_{1}$ (18)
where the explicit definitions of the components are given as follows
$\overset{(2)}{T}{}^{a}$ $=\frac{1}{2}e^{a}\wedge T$$\Leftrightarrow$ $T$ $=\iota_{a}\overset{(2)}{T}{}^{a},$ (19a)
$\overset{(3)}{T}{}^{a}$ $=\frac{1}{3}\iota^{a}\sigma$$\Leftrightarrow$ $\sigma$ $=e_{a}\wedge\overset{(3)}{T}{}^{a},$ (19b)
$\overset{(1)}{T}{}^{a}$ $=T^{a}-\overset{(2)}{T}{}^{a}-\overset{(3)}{T}{}^{a}.$ (19c)
The number under a torsion quantity denotes the number of its components. Since any one-form $T$ and any three-form $\sigma$ have three and one components, respectively, we firstly count those numbers for $\overset{(2)}{T}{}^{a}$ and $\overset{(3)}{T}{}^{a}$. Then we count the components of $T^{a}$ as nine. Finally we compute the components for $\overset{(1)}{T}{}^{a}$ as five. Consequently these components satisfy the properties
$\iota_{a}\overset{(1)}{T}{}^{a}=\iota_{a}\overset{(3)}{T}{}^{a}=0\quad\text{and}\quad e_{a}\wedge\overset{(1)}{T}{}^{a}=e_{a}\wedge\overset{(2)}{T}{}^{a}=0$ (20)
and they are orthogonal in the following sense
$\overset{(i)}{T}{}^{a}\wedge*\overset{(j)}{T}{}_{a}=\delta^{ij}N_{ij}*1,\quad i,j=1,2,3,\qquad\text{(no summation over $ij$)}$ (21)
where $\delta^{ij}$ is Kronecker delta and $N_{ij}$ is a zero-form.
[…]
We use them in the decomposition of non-metricity as
$\underbrace{Q_{ab}}_{18}=\underbrace{\overset{(1)}{Q}{}_{ab}}_{3}+\underbrace{\overset{(2)}{Q}{}_{ab}}_{9}+\underbrace{\overset{(3)}{Q}{}_{ab}}_{3}+\underbrace{\overset{(4)}{Q}{}_{ab}}_{3}$ (24)
where
$\overset{(2)}{Q}{}_{ab}$ $=\frac{1}{3}{}*(e_{a}\wedge\Omega_{b}+e_{b}\wedge\Omega_{a})$$\Leftrightarrow$ $\Theta_{a}$ $=*\left(\overset{(2)}{Q}{}_{ab}\wedge e^{b}\right)-\frac{1}{2}\Lambda_{b}*(e^{b}\wedge e_{a}),$ (25a)
$\overset{(3)}{Q}{}_{ab}$ $=\frac{3}{10}\left(\Lambda_{a}e_{b}+\Lambda_{b}e_{a}-\frac{2}{3}\delta_{ab}\Lambda\right)$$\Leftrightarrow$ $\Lambda_{a}$ $=\iota^{b}\overset{(3)}{Q}{}_{ab},$ (25b)
$\overset{(4)}{Q}{}_{ab}$ $=\frac{1}{3}\delta_{ab}Q$$\Leftrightarrow$ $Q$ $=\delta^{ab}\overset{(4)}{Q}{}_{ab},$ (25c)
$\overset{(1)}{Q}{}_{ab}$ $=Q_{ab}-\overset{(2)}{Q}{}_{ab}-\overset{(3)}{Q}{}_{ab}-\overset{(4)}{Q}{}_{ab}.$ (25d)
We firstly count the components of $\overset{(2)}{Q}{}_{ab}$ as nine, because any one-form $\Theta_{a}$ has nine component. Similarly since a zero-form $\Lambda_{a}$ has three components, $\overset{(3)}{Q}{}_{ab}$ does the same. Then we count it for $\overset{(4)}{Q}{}_{ab}$ as three thanks to any one-form $Q$ has three components. Finally they all must be added up to eighteen for $Q_{ab}$, $\overset{(1)}{Q}{}_{ab}$ must have three components. They satisfy the properties
$\delta^{ab}\overset{(1)}{Q}{}_{ab}=\delta^{ab}\overset{(2)}{Q}{}_{ab}=\delta^{ab}\overset{(3)}{Q}{}_{ab}=0,$ (26a)
$\iota^{a}\overset{(1)}{Q}{}_{ab}=\iota^{a}\overset{(2)}{Q}{}_{ab}=0,$ (26b)
$e^{a}\wedge\overset{(1)}{Q}{}_{ab}=0.$ (26c)
Thus the components are orthogonal in the following sense
$\overset{(i)}{Q}{}^{ab}\wedge\;*\overset{(j)}{Q}{}_{ab}=\delta^{ij}N_{ij}*1,\quad i,j=1,2,3,4,\qquad\text{(no summation over $ij$)}$ (27)
where $\delta^{ij}$ is the Kronecker symbol and $N_{ij}$ is a zero-form. Since we will work in the general teleparallel geometries in which the full curvature vanishes we do not review the decomposition of it. One can read on it in Ref [22].
2.4 Path dependence of scalar product in a non-metric geometry

### v2

# Reformulation of continuum defects in terms of the general teleparallel geometry in the language of exterior algebra

@@ Abstract
We discuss the geometric formulation of continuum defects consisting of dislocations and disclinations. After reviewing the metric affine geometry and the present geometric formulation of dislocation and disclination written in terms of torsion and full curvature (together with vanishing non-metricity), we give a new formulation of them in a novel way in terms of torsion and non-metricity (together with vanishing full curvature), the so-called general teleparallel geometry. All calculations are performed by using the exterior algebra. We obtain continuity equations explicitly for dislocation density and disclination density.
@@ 1 Introduction
It is known that the role of geometry in physics is very valuable. The most familiar example of it is general relativity, Einstein’s theory of gravity, in which the gravitational field is associated with the Riemann curvature tensor. With similar motivations, geometric quantities have been used to express the theory of defects in materials. In this context, the physical state of a material describes the geometry of the space occupied by its atoms. In studies conducted for this purpose so far, dislocation type defects have been paired with torsion and disclination type defects with full curvature [1]-[11]. We review main results of those works in Section 3 in terms of exterior forms. There are works discussing the point defects apart from these two line defects as well and their possible relations to the non-metricity [12]-[17]. In addition to these works, the authors of Ref [18] investigate the nonlinear equations of the elasticity theory for bodies with impurities, dislocations and disclinations by adhering a novel mathematical perspective, more specifically by using the method of the Poisson “hydrodynamic” brackets. The papers [19],[20] may be seen as the first works discussing the possible relations among the non-metric connections (i.e. non-metricity), quasi-dislocations and quasi-disclinations.
@@
However, if torsion is paired with dislocation and full curvature with disclination as in the geometric picture we summarized above, there will be no shift in dislocation when disclination occurs. On the contrary, the presence of dislocation causes a shift in disclination as shown in the papers [21]-[24]. This contradiction is the first motivation leading us to do this work.
On the other hand, we know that a metric affine geometry is defined by the triple; manifold, metric, affine connection or correspondingly by the triple; non-metricity, torsion, curvature. Furthermore, the affine connection 1-form, $\omega_{ab}$, contains a Riemannian component (the so-called Levi-Civita connection 1-form, $\widetilde{\omega}_{ab}$, computed from metric) and a non-Riemannian component (which we call as defect tensor-valued 1-form, $\mathrm{L}_{ab}$, consisting of tensors of torsion and non-metricity). Although in the literature of geometry and gravity it is usually called as distortion 1-form, we prefer to name it as defect 1-form appropriate to the topic. We summarize the basics of metric affine geometry in the forthcoming section. More specifically, while the symmetric part of defect 1-form consist of only non-metricity 1-form, its anti-symmetric part contains torsion and certain components of non-metricity. Thus, when the non-metricity vanishes, the defect 1-form turns out to be anti-symmetric and is given only by torsion. However, when non-metricity does not vanish, along with that the symmetric part of the defect 1-form appears, its anti-symmetric part is shifted by certain components of non-metricity. One can consult Ref.[28] for further reading. In conclusion, the resemblances between the statements on dislocation (disclination) and the anti-symmetric part (symmetric part) of defect 1-form led us to investigate the reformulation of the defect theory of a continuum in terms of Cartan-Weyl (general teleparallel) geometry. Besides all these, on the geometry side, the general teleparallel geometry has an advantage over the Riemann-Cartan geometry. While the latter does not have a metric formulation, the former does. Metric formulation of general teleparallel geometry is the second motivation leading us to do this work. In fact, all three teleparallel geometries have metric formulations [28]. As far as we know, our general teleparallel geometric formulation of continuum defects is the first attempt in the literature. Free from special discussion here, although plenty of papers has been published, no fundamental theory of defects is yet available. We develop our formulation for any non-Cartesian coordinates covariantly in the language of exterior algebra.
@@ 2 Metric affine geometry
The triple $\{M,g,\nabla\}$ or $\{Q_{ab},T^{a},R^{a}{}_{b}\}$ defines a metric affine geometry where $M$ is three-dimensional orientable and differentiable manifold, $g$ is non-degenerate symmetric metric, $\nabla$ is full (or affine) connection or covariant derivative, $Q_{ab}$ is non-metricity 1-form, $T^{a}$ is torsion 2-form and $R^{a}{}_{b}$ is curvature 2-form [29], [30]. We denote the anholonomic metric-orthonormal coframe by $e^{a}$, then write the metric as $g=g_{ab}e^{a}\otimes e^{b}$ where $g_{ab}:=\delta_{ab}$ is the Euclid metric with $\delta_{ab}=\text{diag}(1,1,1)$. The full connection is determined by the full connection 1-form $\omega^{a}{}_{b}$ via the definition $\nabla e^{a}:=-\omega^{a}{}_{b}\wedge e^{b}$ or $\nabla X_{a}:=X_{b}\otimes\omega^{b}{}_{a}$ where $\otimes$ denotes tensor product, $\wedge$ is the exterior product in the exterior algebra and $X_{a}$ is the orthonormal frame such that $e^{a}(X_{b})=\delta^{a}_{b}$ is the duality relation. Besides, $e^{a}$ is called the metric orthonormal (or shortly orthonormal) 1-form and the Cartan structure equations are given by tensor-valued non-metricity 1-form [^1] Here $\omega_{(ab)}=Q_{ab}$ is the reason why we adopt the factor $-\frac{1}{2}$ in the definition of non-metricity, tensor-valued torsion 2-form and tensor-valued full (or non-Riemannian) curvature 2-form, respectively,
$Q_{ab}$ $:=-\frac{1}{2}D\delta_{ab}=-\frac{1}{2}\left(d\delta_{ab}-\omega^{c}{}_{a}\delta_{cb}-\omega^{c}{}_{b}\delta_{ac}\right)=\omega_{(ab)},$ (1a)
@@
where $d$ is the exterior derivative and $D$ is the covariant exterior derivative. We used the result $d\delta_{ab}=0$ in the orthonormal frame in (1a). We utilize the usual notation: ()($\bullet$)( ∙ ) and []delimited-[][$\bullet$][ ∙ ] means to calculate the symmetric and anti-symmetric parts with respect to the indices enclosed, i.e., $\omega_{(ab)}=\frac{1}{2}(\omega_{ab}+\omega_{ba})$ and $\omega_{[ab]}=\frac{1}{2}(\omega_{ab}-\omega_{ba})$. They satisfy the Bianchi identities
@@
$D\mathfrak{T}^{a_{1}a_{2}\cdots a_{p}}_{\;\;\;\;b_{1}b_{2}\cdots b_{q}}$ $=d\mathfrak{T}^{a_{1}a_{2}\cdots a_{p}}_{\;\;\;\;b_{1}b_{2}\cdots b_{q}}+\omega^{a_{1}}{}_{c}\wedge\mathfrak{T}^{ca_{2}\cdots a_{p}}_{\;\;\;\;b_{1}b_{2}\cdots b_{q}}+\cdots+\omega^{a_{p}}{}_{c}\wedge\mathfrak{T}^{a_{1}a_{2}\cdots c}_{\;\;\;\;b_{1}b_{2}\cdots b_{q}}$ (3)
$\quad-\omega^{c}{}_{b_{1}}\wedge\mathfrak{T}^{a_{1}a_{2}\cdots a_{p}}_{\;\;\;\;cb_{2}\cdots b_{q}}-\cdots-\omega^{c}{}_{b_{q}}\wedge\mathfrak{T}^{a_{1}a_{2}\cdots a_{p}}_{\;\;\;\;b_{1}b_{2}\cdots c}.$
@@
Thus, we deduce that $Q_{abc}=Q_{(ab)c}$ is symmetric at the first two indices, $T^{a}{}_{bc}=T^{a}{}_{[bc]}$ is anti-symmetric at the last two indices and $R_{abcd}=R_{ab[cd]}$ is anti-symmetric at the last two indices and asymmetric $R_{abcd}=R_{[ab]cd}+R_{(ab)cd}$ at the first two indices by definition. In the subsequent sections, the definitions of the first kind trace 1-form $Q:=\delta_{ab}Q^{ab}=Q^{a}{}_{ab}e^{b}$ and the second kind trace 1-form $P:=(\iota_{a}Q^{ab})e_{b}=Q^{a}{}_{ba}e^{b}$ of the non-metricity tensor will be used frequently where $\iota_{a}\equiv\iota_{X_{a}}$ denotes the interior derivative of the exterior algebra, $\iota_{b}e^{a}=\delta^{a}_{b}$. Furthermore, it is worthy to remember that the symmetric part of the first two indices of $R_{abcd}$ depends on existence of the non-metricity and will allow us to define metrical disclination density tensor in the next section. Furthermore, we would like to remind explicitly that $d\delta_{ab}=0$, $d\delta^{ab}=0$ and $d\delta^{a}_{b}=0$, but $D\delta_{ab}=-2Q_{ab}$, $D\delta^{ab}=+2Q^{ab}$ and $D\delta^{a}_{b}=0$. Accordingly, one must pay special attention as raising or lowering an index in front of $D$ operation in non-metric geometries.
Meanwhile it is worthwhile to remark shortly on our choice of the orthonormal coframe. In this paper we adhere to formulate all quantities in the orthonormal frame (or equivalently in the orthonormal coframe). Of course, a non-orthogonal coframe can be chosen, for example, the coordinate coframe, $dx^{\alpha}$. Nevertheless, it is known that one can always pass from the orthonormal coframe to the coordinate coframe or vise versa at any stage of calculations via a linear transformation $dx^{\alpha}=h^{\alpha}{}_{a}e^{a}$ and its inverse $e^{a}=h^{a}{}_{\alpha}dx^{\alpha}$ where $h^{\alpha}{}_{a}$ is called as dreibein. Then, the affine connection 1-form transforms correspondingly as $\omega^{\alpha}{}_{\beta}=h^{\alpha}{}_{a}\omega^{a}{}_{b}h^{b}{}_{\beta}+h^{\alpha}{}_{a}dh^{a}{}_{\beta}$. Thus, tensor-valued exterior forms transform covariantly, i.e., $g_{\alpha\beta}=h^{a}{}_{\alpha}h^{b}{}_{\beta}\delta_{ab}$, $Q_{\alpha\beta}=h^{a}{}_{\alpha}h^{b}{}_{\beta}Q_{ab}$, $T^{\alpha}=h^{\alpha}{}_{a}T^{a}$, $R^{\alpha}{}_{\beta}=h^{\alpha}{}_{a}R^{a}{}_{b}h^{b}{}_{\beta}$. Finally we rewrite the Cartan structure equations in the coordinate coframe as
$Q_{\alpha\beta}$ $:=-\frac{1}{2}Dg_{\alpha\beta}=-\frac{1}{2}\left(dg_{\alpha\beta}-\omega^{\gamma}{}_{\alpha}g_{\gamma\beta}-\omega^{\gamma}{}_{\beta}g_{\alpha\gamma}\right)=-\frac{1}{2}dg_{\alpha\beta}+\omega_{(\alpha\beta)},$ (5a)
$T^{\alpha}$ $:=D(dx^{\alpha})=d(dx^{\alpha})+\omega^{\alpha}{}_{\beta}\wedge dx^{\beta}=\omega^{\alpha}{}_{\beta}\wedge dx^{\beta},$ (5b)
$R^{\alpha}{}_{\beta}$ $:=D\omega^{\alpha}{}_{\beta}:=d\omega^{\alpha}{}_{\beta}+\omega^{\alpha}{}_{\gamma}\wedge\omega^{\gamma}{}_{\beta},$ (5c)
where $d(dx^{\alpha})=d^{2}x^{\alpha}=0$ because of Poincare lemma in the exterior algebra. One of advantages of the usage of the orthonormal coframe is that the exterior derivative and the variation of metric components vanish, $dg_{ab}=0$ and $\delta g_{ab}=0$, but they do not in the non-orthonormal coframe, $dg_{\alpha\beta}\neq 0$ and $\delta g_{\alpha\beta}\neq 0$. For further discussion on the orthonormal coframe, the coordinate coframe and the mixed coframe we direct the reader to the paper [28].
@@ 2.1 Decomposition of full connection
In the literature it is common to define the tensor-valued contortion 1-form, $K_{ab}=-K_{ba}$, in terms of torsion 2-form
@@ 2.2 Some algebraic identities
where * denotes the Hodge duality map, $\Lambda$ is any $p$-form, $\Gamma$ is any $q$-form and $pq$ is the ordinary multiplication of numbers of $p$ and $q$. The operation of Hodge map may be performed via the totally anti-symmetric epsilon tensor by relations
@@
2.3 Path independence of scalar product in the non-metric teleparallel geometries